OpenStudy (anonymous):

Solve log(x+1)- log(x-1)=1

OpenStudy (mertsj):

log((x+1/(x-1)]=1 10=(x+10/(x-1) 10(x-1)=x+1 10x-10=x+1 9x=11 x=11/9